# J-CHAT: Japanese Large-scale Spoken Dialogue Corpus for Spoken Dialogue Language Modeling

### v1

@@ 5 Experiments
5.1 Experimental Conditions
We split the dialogues in the J-CHAT corpus into two channels by swapping the output channel of speech according to the turn-taking event. Subsequently we randomly split the each subset of J-CHAT corpus into train/valid/test sets. Then, we performed discretization of speech using k-means clustering on HuBERT-extracted features using train set from both subsets. The number of clusters for k-means was set to 1000.
For the vocoder used for speech generation from the discretized speech, we used HiFi-GAN Kong et al. (2020) conditioned with speaker information from XVector Snyder et al. (2018), as used in previous work Kharitonov et al. (2022). Please refer to appendix A for more details.
5.2 Evaluation
For evaluation, we performed subjective listening tests on the mean opinion score (MOS) on the naturalness and meaningfulness of dialogues following previous work Nguyen et al. (2023). For each subjective listening test, we recruited 60 Japanese native speakers, and each listener evaluated 8 samples. For sample generation used in the evaluation, we used the first 5 seconds of the test set derived from J-CHAT to prompt the model, then performed inference to predict the next 25 seconds of the dialogue based on those initial 5 seconds. For the sampling method, we used the beam search with a beam size of 5. When synthesizing speech from discritized speech, we conditioned the vocoder with JVS001 (male) and JVS002 (female) speaker from the JVS corpus Takamichi et al. (2020). This resulted in the male speaker’s voice being heard from the first audio channel and the female speaker’s voice from the second channel.
5.3 Result
Table 3: Result for MOS tests with their 95% confidence intervals.
@@ 5.3 Result
Table 3 shows the results of the subjective evaluation. From these results, it can be seen that dGSLM-J-CHAT achieves the best performance among the dGSLM generated samples in both naturalness and meaningfulness. This suggests that J-CHAT is a useful corpus for constructing generative dialogue language models. We can also see that there is no statistical significance between dGSLM-YouTube and dGSLM-podcast, despite dGSLM-podcast being trained on a dataset approximately four times larger in number of dialogues. This indicates that simply scaling up the dataset is not enough to enhance dGSLM performance. However, there is a significant difference between resynth and dGSLM-J-CHAT in terms of both naturalness and meaningfulness. The trained model occasionally produces sensible words, but the generated dialogue often lacks coherence. This might be improved with better modeling or by increasing the dataset size. For a detailed explanation on the statistical significance of each result, refer to the appendix B.2.
6 Conclusion
We have demonstrated a scalable dataset construction method. The experimental results showed that collecting from multiple domains, such as YouTube and podcasts, is beneficial in training dialogue language models. We hope that this corpus will contribute to the advancement of research in dialogue speech models.
Limitations
In this study, we believe that obtaining data from multiple domains, such as YouTube and podcasts, allows us to gather a wide range of speech. However, we also recognize that there are limitations to the range that can be covered. For example, OGVC corpus Arimoto et al. (2012) reports a method to collect a laughter corpus from the internet, but found that the amount of screaming was limited because the internet data is typically collected under controlled conditions. Similarly, podcast dialogues are intended for listeners, so informal expressions used among close friends are likely to be limited.
Ethics Statement
We recognize that speech synthesis technology carries the risk of being misused for unauthorized generation of others’ voices, commonly referred to as voice cloning. To prevent such misuse, we converted the collected speech into latent variables known as HuBERT features, which lack speaker identity information. Although HuBERT does not completely eliminate speaker information, it obscures it sufficiently to make the restoration of speaker identity challenging, thus serving as an effective countermeasure against voice cloning. In our experiments, we synthesized speech using the speaker identity of individuals who had consented to their voice being used for synthesis at the time of recording, ensuring that no unauthorized voice replication was performed.
On the subjective evaluation conducted on Section 5.2, participants are paid with adequate amount of money and they were instructed with the purpose of the research and how the collected data will be used. Please refer to the Section B for the details on the subjective evaluation.
Acknowledgements
@@
Appendix A dGSLM Training Details
For the implementation of dGSLM, we used fairseq [^3] https://github.com/facebookresearch/fairseq repository by Meta which includes official implementation of the model. This repository is licensed as MIT.
The number of train/valid/test is 1,003,397/9,991/100 dialogues for YouTube subset and for podcast subset it is 3,885,007/38,902/100 dialogues.
For HuBERT model used in the speech-to-unit, we used the pretrained Japanese HuBERT model Sawada et al. (2024) [^4] https://huggingface.co/rinna/japanese-hubert-base. For k-means clustering, we used the original implementation of the dGSLM except for the number of clusters which is set to 1,000.
The dGSLM model was configured as in the previous research Nguyen et al. (2023). The number of model parameters in each dGSLM model is 23,638,529.
For training dGSLM model, we used 32 NVIDIA V100 GPUs with the learning rate of $2\times 10^{-4}$. The training was performed for 100,000 steps. The total duration of training was around 48 hours. The batch size was 36,864 tokens and the the maximum number of tokens per sample are set to 3,000. This is equivalent to the 1 minute of dialogue. For other hyper-parameters regarding the training, we followed the original paper Nguyen et al. (2023).
For the XVector conditioned vocoder, we used the pretrained XVector model [^5] https://huggingface.co/speechbrain/spkrec-xvect-voxceleb. The vocoder was trained with the JVS Takamichi et al. (2020) and JVNV Xin et al. (2024a) corpus. These two corpora include Japanese reading-style, studio-quality speech without/with non-verbal expression, respectively. Training of vocoder took 2 days with 4 NVIDIA V100 GPUs.
Appendix B Subjective Evaluation of the dGSLM Outputs
B.1 Participant Details
Participants of the each subjective evaluation were recruited on https://lancers.jp. They are paid with 75 Japanese yen upon the completion of the test. Given that the minimum wage of Japan is 1,004 Japanese yen and the test took approximately 4 minutes, the payment is adequate.
Before the participation, the participants were instructed with the research purpose and how the collected data will be used in this research. The data collection protocol is approved by our institution’s ethics review board.
Figure 3: Screenshot of instruction given to the participants on the naturalness MOS evaluation.
Figure 4: Screenshot of instruction given to the participants on the meaningfulness MOS evaluation.
Figures  3 and 4 show the screenshots of the instructions given to the participants for the naturalness and meaningfulness MOS evaluations, respectively. The English translation of the instruction for the naturalness MOS evaluation is as follows:
Listen to the audio and rate the naturalness of the sound on a scale of 1 to 5 (very unnatural (1) to very natural (5)). Please conduct the evaluation using headphones in a quiet environment. Please note that some audio samples may have degraded sound quality, so please be careful with the volume level. Do not use your browser’s “back” or “reload” buttons during the test.
The English translation of the instruction for the meaningfulness MOS evaluation is as follows:
Listen to the audio and rate the meaningfulness of the dialogue on a 5-point scale from very meaningless (1) to very meaningful (5). Please evaluate in a quiet environment using headphones. Please note that some of the audio samples may have degraded sound quality, so please be careful with the volume. Please do not use your browser’s “back” or “reload” buttons during the test.
B.2 P-values from the Subjective Evaluation
Figure 5: P-values from the subjective evaluation on the naturalness of the dGSLM output samples
Figure 6: P-values from the subjective evaluation on the meaningfulness of the dGSLM output samples
Figure 5 and Figure 6 shows the p-values from the subjective evaluation on the dGSLM output conducted on Section 5.2. Specifically, we performed two sided T-test on the MOS score.

### v2

@@ 5.  Experiments
5.1.  Experimental Conditions
We split the dialogues in the J-CHAT corpus into two channels by swapping the output channel of speech according to the turn-taking event. We used the train/valid/test sets explained in Section 4.1. Then, we performed discretization of speech using k-means clustering on HuBERT-extracted features using train set from both YouTube and Podcast subsets. The number of clusters for k-means was set to 1,000.
For the vocoder used for speech generation from the discretized speech, we used HiFi-GAN Kong et al. (2020) conditioned with speaker information from XVector Snyder et al. (2018), as used in previous work Kharitonov et al. (2022).
5.2.  Training Details
For the implementation of dGSLM, we used official implementation of the model [^6] https://github.com/facebookresearch/fairseq. For HuBERT model used in the speech-to-unit, we used the pretrained Japanese HuBERT model Sawada et al. (2024) [^7] https://huggingface.co/rinna/japanese-hubert-base. For k-means clustering, we used the original implementation of the dGSLM except for the number of clusters which was set to $1,000$. The dGSLM model was configured as in the previous research Nguyen et al. (2023).
For training dGSLM model, we used 32 NVIDIA V100 GPUs with the learning rate of $2\times 10^{-4}$. The training was performed for $100,000$ steps. The batch size was $36,864$ tokens and the the maximum number of tokens per sample are set to $3,000$. This was equivalent to the $1$ minute of dialogue. For other hyper-parameters regarding the training, we followed the original paper Nguyen et al. (2023).
Table 3: MOS test results with their 95% confidence intervals.
@@ 5.2.  Training Details
For the XVector conditioned vocoder, we used the pretrained XVector model [^8] https://huggingface.co/speechbrain/spkrec-xvect-voxceleb. The vocoder was trained with the JVS Takamichi et al. (2020) and JVNV Xin et al. (2024) corpus. These two corpora include Japanese reading-style, studio-quality speech without/with non-verbal expression, respectively. Training of vocoder took 2 days with 4 NVIDIA V100 GPUs.
5.3.  Evaluation
For evaluation, we performed subjective listening tests on the mean opinion score (MOS) on the naturalness and meaningfulness of dialogues following previous work Nguyen et al. (2023). For each subjective listening test, we recruited 60 Japanese native speakers, and each listener evaluated 8 samples. For sample generation used in the evaluation, we used the first 5 seconds of the test set derived from J-CHAT to prompt the model, then performed inference to predict the next 25 seconds of the dialogue based on those initial 5 seconds. For the sampling method, we used the beam search with a beam size of 5. When synthesizing speech from discritized speech, we conditioned the vocoder with JVS001 (male) and JVS002 (female) speaker from the JVS corpus Takamichi et al. (2020). This resulted in the male speaker’s voice being heard from the first audio channel (left) and the female speaker’s voice from the second channel (right).
5.4.  Result
@@
Table 3 shows the results of the subjective evaluation. From these results, it can be seen that dGSLM-J-CHAT achieves the best performance among the dGSLM generated samples in both naturalness and meaningfulness. This suggests that J-CHAT is a useful corpus for constructing generative dialogue language models. We can also see that there is no statistical significance between dGSLM-YouTube and dGSLM-podcast, despite dGSLM-podcast being trained on a dataset approximately four times larger in number of dialogues. This indicates that simply scaling up the dataset is not enough to enhance dGSLM performance. However, there is a significant difference between resynth and dGSLM-J-CHAT in terms of both naturalness and meaningfulness. The trained model occasionally produces sensible words, but the generated dialogue often lacks coherence. This might be improved with better modeling or by increasing the dataset size.
6.  Conclusion
In this study, we presented J-CHAT, a large-scale Japanese spoken dialogue corpus, and proposed an automated and language-independent methodology for its construction. The experimental results demonstrated that training on diverse dialogue data from multiple data sources, such as YouTube and podcasts, significantly improves spoken dialogue systems models than collecting from single data source. Future work includes further improving dialogue modeling techniques to enhance the quality of generated speech.
7.  Ethics statement
The published data were collected and distributed in Japan and therefore comply with Japanese law. The legality of the collection and distribution process of J-CHAT was reviewed by a third-party Japanese lawyer. Under Article 30-4 of the Japanese Copyright Act, copyrighted materials may be used without explicit permission when the purpose is limited to “information analysis.”2 Accordingly, the dataset is released strictly for purposes related to information analysis, and not for entertainment or content redistribution. To mitigate privacy concerns, we performed anonymization of source metadata. In addition, the original recordings were segmented into dialogue units, preventing reconstruction of the original content for entertainment purposes. We also implement an opt-out policy, following prior large-scale web corpus practices (e.g. Common Crawl), allowing individuals or rights holders to request removal of their data from the dataset.
8.  Acknowledgements
@@
9.  Bibliographical References
  • V. Agrawal, A. Akinyemi, K. Alvero, M. Behrooz, J. Buffalini, F. M. Carlucci, J. Chen, J. Chen, Z. Chen, S. Cheng, et al. (2025) Seamless interaction: dyadic audiovisual motion modeling and large-scale dataset. arXiv preprint arXiv:2506.22554. Cited by: §1, §1, §2.
  • C. Alexandra and G. Zipperlen (1996) CALLHOME Japanese speech LDC96S37. In Linguistic Data Consortium, Cited by: §1, Figure 2, §4.3.
  • Z. Borsos, M. Sharifi, D. Vincent, E. Kharitonov, N. Zeghidour, and M. Tagliasacchi (2023) SoundStorm: efficient parallel audio generation. External Links: 2305.09636 Cited by: §1.
  • G. Chen, S. Chai, G. Wang, J. Du, W. Zhang, C. Weng, D. Su, D. Povey, J. Trmal, J. Zhang, M. Jin, S. Khudanpur, S. Watanabe, S. Zhao, W. Zou, X. Li, X. Yao, Y. Wang, Z. You, and Z. Yan (2021) GigaSpeech: An Evolving, Multi-Domain ASR Corpus with 10,000 Hours of Transcribed Audio. In Proc. Interspeech 2021, pp. 3670–3674. External Links: Document, ISSN 2958-1796 Cited by: §1, §1, §2.
  • C. Cieri, D. Miller, and K. Walker (2004) The Fisher Corpus: a resource for the next generations of speech-to-text. In Proceedings of the Fourth International Conference on Language Resources and Evaluation (LREC’04), M. T. Lino, M. F. Xavier, F. Ferreira, R. Costa, and R. Silva (Eds.), Lisbon, Portugal. External Links: Link Cited by: §1, §2.
  • A. Défossez, L. Mazaré, M. Orsini, A. Royer, P. Pérez, H. Jégou, E. Grave, and N. Zeghidour (2024) Moshi: a speech-text foundation model for real-time dialogue. arXiv arXiv:2410.00037. Cited by: §3.4.
  • W. Hsu, B. Bolte, Y. H. Tsai, K. Lakhotia, R. Salakhutdinov, and A. Mohamed (2021) HuBERT: self-supervised speech representation learning by masked prediction of hidden units. IEEE/ACM Transactions on Audio, Speech, and Language Processing 29 (), pp. 3451–3460. External Links: Document Cited by: Figure 3, §4.3.
  • R. Huang, M. Li, D. Yang, J. Shi, X. Chang, Z. Ye, Y. Wu, Z. Hong, J. Huang, J. Liu, et al. (2024) Audiogpt: understanding and generating speech, music, sound, and talking head. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38, pp. 23802–23804. Cited by: §1.
  • E. Kharitonov, A. Lee, A. Polyak, Y. Adi, J. Copet, K. Lakhotia, T. A. Nguyen, M. Riviere, A. Mohamed, E. Dupoux, and W. Hsu (2022) Text-free prosody-aware generative spoken language modeling. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), S. Muresan, P. Nakov, and A. Villavicencio (Eds.), Dublin, Ireland, pp. 8666–8681. External Links: Link, Document Cited by: §5.1.
  • Y. Koizumi, H. Zen, S. Karita, Y. Ding, K. Yatabe, N. Morioka, M. Bacchiani, Y. Zhang, W. Han, and A. Bapna (2023) LibriTTS-R: a restored multi-speaker text-to-speech corpus. In Proc. INTERSPEECH 2023, pp. 5496–5500. External Links: Document, ISSN 2958-1796 Cited by: §3.3.
  • J. Kong, J. Kim, and J. Bae (2020) HiFi-GAN: generative adversarial networks for efficient and high fidelity speech synthesis. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M.F. Balcan, and H. Lin (Eds.), Vol. 33, pp. 17022–17033. External Links: Link Cited by: §5.1.
  • K. Lakhotia, E. Kharitonov, W. Hsu, Y. Adi, A. Polyak, B. Bolte, T. Nguyen, J. Copet, A. Baevski, A. Mohamed, et al. (2021) On generative spoken language modeling from raw audio. Transactions of the Association for Computational Linguistics 9, pp. 1336–1354. External Links: Link, Document Cited by: §5.
  • K. Lee, K. Park, and D. Kim (2023) DailyTalk: spoken dialogue dataset for conversational text-to-speech. In ICASSP 2023 - 2023 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), Vol. , pp. 1–5. External Links: Document Cited by: §1.
  • K. Mitsui, Y. Hono, and K. Sawada (2023) Towards human-like spoken dialogue generation between ai agents from written dialogue. External Links: 2310.01088 Cited by: §1.
  • G. Mittag, B. Naderi, A. Chehadi, and S. Möller (2021) NISQA: A Deep CNN-Self-Attention Model for Multidimensional Speech Quality Prediction with Crowdsourced Datasets. In Proc. Interspeech 2021, pp. 2127–2131. External Links: Document, ISSN 2958-1796 Cited by: §4.2.
  • T. A. Nguyen, E. Kharitonov, J. Copet, Y. Adi, W. Hsu, A. Elkahky, P. Tomasello, R. Algayres, B. Sagot, A. Mohamed, and E. Dupoux (2023) Generative spoken dialogue language modeling. Transactions of the Association for Computational Linguistics 11, pp. 250–266. External Links: Link, Document Cited by: §1, §1, §5.2, §5.2, §5.3, §5.
  • S. Park, C. Kim, H. Rha, M. Kim, J. Hong, J. Yeo, and Y. Ro (2024) Let’s go real talk: spoken dialogue model for face-to-face conversation. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), L. Ku, A. Martins, and V. Srikumar (Eds.), pp. 16334–16348. External Links: Link, Document Cited by: §1, §2.
  • A. Plaquet and H. Bredin (2023) Powerset multi-class cross entropy loss for neural speaker diarization. In Proc. INTERSPEECH 2023, pp. 3222–3226. External Links: Document, ISSN 2958-1796 Cited by: §3.2.2.
  • S. Rouard, F. Massa, and A. Défossez (2023) Hybrid transformers for music source separation. In ICASSP 2023 - 2023 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), Vol. , pp. 1–5. External Links: Document Cited by: §3.3.
  • Y. Saito, Y. Nishimura, S. Takamichi, K. Tachibana, and H. Saruwatari (2022) STUDIES: Corpus of Japanese Empathetic Dialogue Speech Towards Friendly Voice Agent. In Proc. Interspeech 2022, pp. 5155–5159. External Links: Document, ISSN 2958-1796 Cited by: §1, §2, Figure 2, §4.3.
  • K. Sawada, T. Zhao, M. Shing, K. Mitsui, A. Kaga, Y. Hono, T. Wakatsuki, and K. Mitsuda (2024) Release of pre-trained models for the Japanese language. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), N. Calzolari, M. Kan, V. Hoste, A. Lenci, S. Sakti, and N. Xue (Eds.), Torino, Italia, pp. 13898–13905. External Links: Link Cited by: §5.2.
  • K. Seki, S. Takamichi, T. Saeki, and H. Saruwatari (2023) TEXT-to-speech synthesis from dark data with evaluation-in-the-loop data selection. In ICASSP 2023 - 2023 IEEE International Conference on Acoustics, Speech and Signal Processing, Vol. , pp. 1–5. External Links: Document Cited by: §1.
  • Z. Sheikh, S. Shimizu, S. Arora, J. Shi, S. Cornell, X. Li, and S. Watanabe (2025) Scalable spontaneous speech dataset (SSSD): crowdsourcing data collection to promote dialogue research. In Proc. Interspeech 2025, pp. 3963–3967. Cited by: §1, §2.
  • D. Snyder, D. Garcia-Romero, G. Sell, D. Povey, and S. Khudanpur (2018) X-Vectors: robust dnn embeddings for speaker recognition. In 2018 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), Vol. , pp. 5329–5333. External Links: Document Cited by: §5.1.
  • S. Takamichi, L. Kürzinger, T. Saeki, S. Shiota, and S. Watanabe (2021) JTubeSpeech: corpus of japanese speech collected from youtube for speech recognition and speaker verification. External Links: 2112.09323 Cited by: §1, §3.1.
  • S. Takamichi, R. Sonobe, K. Mitsui, Y. Saito, T. Koriyama, N. Tanji, and H. Saruwatari (2020) JSUT and JVS: free japanese voice corpora for accelerating speech synthesis research. Acoustical Science and Technology 41 (5), pp. 761–768. External Links: Document Cited by: §5.2, §5.3.
  • D. Xin, J. Jiang, S. Takamichi, Y. Saito, A. Aizawa, and H. Saruwatari (2024) JVNV: a corpus of japanese emotional speech with verbal content and nonverbal expressions. IEEE Access 12 (), pp. 19752–19764. External Links: Document Cited by: §5.2.
  • H. Zen, V. Dang, R. Clark, Y. Zhang, R. J. Weiss, Y. Jia, Z. Chen, and Y. Wu (2019) LibriTTS: a corpus derived from LibriSpeech for text-to-speech. arXiv preprint arXiv:1904.02882. Cited by: §1, §2.